Accelerator Update for November 3, 2008 - November 5, 2008
Monday Nov 3
The day shift began with the Operations monitoring the Tevatron (TeV) store 6539 (112.42E30) and the Recycler stash (194.0E10), with the Antiproton Source (Pbar) stacking (26.5E10), with the Linac, Booster, and the Main Injector (MI) operating normally, with the MiniBooNE and NuMI experiments taking beam, and with the Collider Detector at Fermilab (CDF) and D-Zero (D0) using the colliding beams.
At 11:12 AM, Operations received a building alarm for the TeV sector F3 service building. Operations also reported a problem with a Switchyard (SY) power supply (HP3US), which was located at F3. Operators responded, along with the Fire Department, Security, and Electrical support technicians. The experts discovered that a capacitor had shorted and then leaked.
Operations terminated store 6539 at 3:34 PM.
Operations put the TeV into shot setup at 3:48 PM.
Operations reported at 4:01 PM that the Recycler had stashed antiprotons seven times during the day shift.
At 5:09 PM, Operations reported the TeV sector D0 cryo system compressor #1 had tripped off. The cryo system techs at CHL tried to start compressor #2, but it also tripped off. Operations contacted the Cryo Coordinator.
At 5:51 PM, a D0 expert reported that their cryo compressor was just brought back online. He reported that it would take some time for their solenoid and central tracker to be ready for physics.
Operations established store 6540 at 5:56 PM with an initial luminosity, reported only by CDF, of 315.1E30. Later, it was estimated that the average luminosity was somewhere between 307 and 310E30.
The Cryo Coordinator reported at 8:54 PM that since the D0 compressor continued to trip they were going to hook the D0 cryo system to the TeV.
Operations reported at midnight that the Recycler had stashed six times during the evening shift.
Tuesday Nov 4
The midnight shift began with Operations monitoring store 6540 (131.98E30) and the stash (191.4E10), with Pbar stacking (22.9E10), and with the MiniBooNE and NuMI experiments taking beam.
At 12:09 AM, Operations reported losing most of the antiprotons during a transfer to the Recycler. The next transfer went okay.
Operations put Pbar into standby at 6:38 AM.
Operations reported at 7:48 AM that the Recycler had stashed seven times during the midnight shift.
Operations reported at 9:22 AM that two SY power supplies (HP3DS and HP3US) had been repaired and they were ready to send beam to Meson.
TeV experts began an end-of-store study at 11:09 AM.
At 12:02 PM, KRF1 tripped off on a LCW flow fault. It reset.
TeV experts completed their study at 1:05 PM.
Operations terminated store 6540 at 1:10 PM.
Operations put the TeV into shot setup at 1:23 PM.
An expert reported at 3:04 PM that he was beginning a Recycler kicker study. At the same time a controls expert reported that he would conduct a troubleshooting study of the Front End (FE) for the Electron Cooling (ECOOL) VME crate.
Operations established store 6541 at 3:06 PM with an initial luminosity of 332.12E30. This store set a New Record Luminosity.
Operations reported at 3:59 PM that the Recycler had stashed four times during the day shift.
The Recycler kicker study ended at 4:21 PM and the ECOOL FE troubleshooting also ended.
At 10:04 PM, LRF3 and LRF4 tripped off on a LCW interlock indication. Operators investigated and then called in Linac experts.
Operations reported at 10:15 PM that they had put Pbar into standby mode and loaded an energy conservation timeline into the Time Line Generator (TLG).
Linac experts reported that one of the L2 transformer breaker had opened. The Duty Electrician reset the trip. Operations began the Linac turn on at 11:40 PM.
Operations reported at 11:57 PM that the Recycler had stashed six times during the evening shift.
Wednesday Nov 5
The midnight shift began with Operations monitoring store 6541 (107.08E30) and the stash (178.8E10), with Pbar stacking (23.2E10), and with the MiniBooNE and NuMI experiments taking beam.
At 1:16 AM, LRF4 tripped off.
The Duty Electrician reset the LRF4 breaker at 2:07 AM.
Operations reported at 2:51 AM that Linac was up and running and sending beam, but that there were MI-10 kicker problems.
Operations established beam to MiniBooNE at 3:13 AM.
At 3:30 AM, the LRF4 breaker tripped off again.
A MI kicker expert reported at 5:41 AM that he had replaced a bad fuse on the MI-10 kicker.
Operations prepared the Booster tunnel for interlock at 6:14 AM. Booster techs will replace the power amplifier for BRF11.
A LCW expert began replacing the flow switch for the LRF2 water skid at 7:11 AM.
Operations turned off the Pelletron at 7:54 AM so a controls expert could replace a FE.
The Plans for Wednesday
The plans for today are to bring Linac back online and resume stacking, stashing, and storing.
Comments and Suggestions
|last modified 11/05/2008 email Fermilab|
|Security, Privacy, Legal|